# Jules Drean MIT CSAIL Miguel Gomez-Garcia MIT CSAIL Fisher Jepsen MIT CSAIL Thomas Bourgeat EPFL  and  Srinivas Devadas MIT CSAIL

### v3

@@ Abstract.
Enclaves or Trusted Execution Environments are trusted-hardware primitives that make it possible to isolate and protect a sensitive program from an untrusted operating system. Unfortunately, almost all existing enclave platforms are vulnerable to microarchitectural side channels and transient execution attacks, and the one academic proposal that is not does not allow programs to interact with the outside world. We present Citadel, to our knowledge, the first enclave platform with microarchitectural isolation to run realistic secure programs on a speculative out-of-order multicore processor. We show how to leverage hardware/software co-design to enable shared memory between an enclave and an untrusted operating system while preventing speculative transmitters between the enclave and a potential adversary. We then evaluate our secure baseline and present further mechanisms to achieve reasonable performance for out-of-the-box programs. Our multicore processor runs on an FPGA and boots untrusted Linux from which users can securely launch and interact with enclaves. To demonstrate our platform capabilities, we run a private inference enclave that embed a small neural network trained on MNIST. A remote user can remotely attest the enclave integrity, perform key exchange and send encrypted input for secure evaluation. We open-source our end-to-end hardware and software infrastructure, hoping to spark more research and bridge the gap between conceptual proposals and FPGA prototypes.
@@ 1. Introduction
Secure Enclaves or Trusted Execution Environments (TEEs) (costan2016sanctum, ; aegis, ; ferraiuolo2017komodo, ; lee2020keystone, ; bahmani2021cure, ; feng2021scalable, ), are trusted-hardware primitives used to protect the execution of sensitive programs. They make it possible to run sensitive computations in isolation with strengthen privacy and integrity properties. Academic TEE proposals date back twenty years to XOM (xom, ) and AEGIS (aegis, ) and industry implementations include ARM TrustZone (trustzone04, ; pinto2019demystifying, ), Intel SGX (mckeen2013sgx, ; costan2016intel, ) and AMD SEV (sev2020strengthening, ).
Unfortunately, these kinds of platforms have been shown to be vulnerable to microarchitectural side channels and transient execution attacks (chen2019sgxpectre, ; koruyeh2018spectre, ; ryan2019hardware, ; van2020lvi, ). In particular, attacks such as those in (van2018foreshadow, ; van2020sgaxe, ) were able to not only extract sensitive data from Intel SGX enclaves, but also managed to break the security of the remote attestation mechanism.
One academic proposal (bourgeat2019mi6, ) has looked at securing enclaves in the context of microarchitectural side channels and transient execution attacks. They aim for a strong security guarantee where a potential attacker can only observe the enclave’s output. Unfortunately, this extreme isolation strongly restricts the class of programs that can be run, as no communication with the outside world is allowed except when entering or exiting the enclave.
@@
We are not the first to observe that restrictive enclave execution models limit the programmability and the performance of enclave programs(van2023cheri, ; yu2022elasticlave, ; li2021confidential, ; lee2022cerberus, ; park2020nested, ; steinegger2021servas, ) but those work ignore security challenges introduced by microarchitectural side-channels and transient execution attacks. When considering those challenges, even basic communication primitives like public shared memory with the operating system are not straightforward to implement efficiently and securely.
In this paper, we try to fill the gap between platform that have strong security but restrict enclave programs to the extreme and platforms that allow better enclave execution models but ignore side channels and transient execution attacks.
We first define our relaxed security goal: an attacker should only be able to observe the transcript of non-speculative accesses to a shared memory region. We then craft requirements for the hardware and the software so that, together, they can prevent secrets from leaking through speculative access to shared memory. In practice, the hardware will isolate private microarchitectural state and restrict the victim program speculation model to help it achieve the just described security guarantee. Intuitively, the more restrictive the speculation model, the lighter the requirements on the program, but the least performance we get.
In our case, we first consider a very conservative speculation model where no speculative shared accesses are allowed. This effectively prevents the existence of any speculative transmitters between the victim and the attacker. We describe how to enable shared memory securely for the specific open-source out-of-order design we used, while disabling speculative shared accesses. We also make sure to not enable any new speculative transmitters by bypassing the L1 D cache, and to not break address translation by tagging the TLB’s and the Address Translation Cache’s entries. Finally, we give a brief overview of the software infrastructure we have implemented in order to boot Linux and run enclaves from user programs on our FPGA prototype.
We use this secure baseline to evaluate the performance of the kind of programs that one could be interested in putting in an enclave. Out-of-the-box, we obtain an underwhelming $120.6\%$ performance overhead for a cryptographic library. We identify and analyze the performance bottlenecks to design a series of hardware-software mechanisms to improve both enclaves’ and unprotected processes’ performance. More precisely, we develop a reconfigurable Last Level Cache (LLC) partitioning scheme, a performance necessity for our small LLC. This reduces our overhead from $120.6\%$ to $19.4\%$. We also introduce software controls to enable a slightly less conservative enclave speculation model, where pipelined accesses to shared memory are allowed but predictors are hardwired to $pc+4$. The efficient way to use this mode is on code segments with burst accesses to shared memory. Under this less conservative speculation model, the programmer needs to make sure that straight-line speculation ($pc+4$) within these segments will not leak any secrets. This mechanism reduces overheads from $19.4\%$ to $2.7\%$ on our cryptographic benchmark.
@@
Our final prototype is a modified out-of-order RISC-V 2-core processor synthesized to run on an AWS F1 FPGA. It fully implements all our security and dynamic resizing mechanisms, boots (untrusted) Linux and makes it possible for user-level applications to launch and safely communicate with secure enclaves executing real applications such as a lightweight Python runtime, a cryptographic library or a private inference application running a small neural network. We also implement an end-to-end attestation mechanism, making it possible for an enclave to identify itself independently of where the OS placed it in physical memory, enabling enclaves to act as a trusted third party. We implement and make all the pieces of software and modified hardware available, such as our secure bootloader, our security monitor and a Linux kernel module to interact with the security monitor. We summarize the different elements implemented for this paper in Figure 1. This makes our platform a unique environment for research in the domain of microarchitectural side channels, transient execution attacks and secure enclaves. We hope researchers will take ownership of the platform, attack it, modify it to prototype and test complex hardware-software mechanisms and research ideas.
@@
  • To our knowledge, we show the first end-to-end enclave platform running securely on an FPGA implementation of a speculative out-of-order processor.
  • We show how to build a shared memory mechanism while guaranteeing an attacker will only observe non-speculative accesses to shared memory.
  • Guided by early performance evaluation, we propose a series of hardware-software mechanisms to make the platform more configurable and improve both enclaves’ and unprotected processes’ performance.
@@
2. Motivation & Background
We want to : 1) run programs that can interact with the outside world. 2) formalize the guarantees that we expect the hardware to provide 3) have good performance.
@@ 2.1. Security Monitor
In many enclave platforms, including Citadel, the Security Monitor (SM) (costan2016sanctum, ) (lebedev2019sanctorum, ) is a small (9K LOC), trusted piece of software running at a higher privilege mode than the hypervisor or the OS. Its role is to link low-level invariants exposed by the hardware (e.g., if a specific memory access is authorized), invoking the right microarchitectural cleanup routine, and the high-level security policies defined by the platform (e.g., which enclave is currently executing on which core and which memory region does it own).
@@ 2.2. Microarchitectural Timing Side Channels
To summarize the observations that could be made through side-channels, the community formalizes side-channels as a set of observations that an attacker could make about the microarchitectural state. For example, cache side-channels on unprotected machines typically lead to an observation model that includes the sequence of addresses of memory operations (sometimes with their timestamps) as part of the observation.
2.3. Transient Execution Attacks and Execution Models
Speculative or transient execution is one of the key performance features of modern processors. Unfortunately, it can make reasoning about security boundaries more difficult. Some microarchitectural data structures that orchestrate speculation (e.g., branch predictors) can be trained so the processor will execute code snippets speculatively, in ways that are incorrect or on data that should not be accessed during normal execution. Concretely, instructions that were not part of the program normal execution path can be executed out-of-thin-air, speculatively, by the processor. An example is the execution of code speculatively before the resolution of a preceding security check, such as a memory access bound. These erroneous executions happen speculatively, and these instructions will be squashed by the processor. Nonetheless, they will still change the microarchitectural state of the machine. For example, a speculative load to some data will not affect the values of the architectural registers but will load the data in the memory hierarchy and possibly change the state of caches. This means that a program that manages sensitive data can be coerced into leaking private data speculatively through a side channel. More importantly, once some specific code snippet, or gadget, has been identified, these attacks can be used to leak virtually any data located in the victim’s virtual address space. This type of attack, called Spectre, was first introduced in (Kocher2018spectre, ), and many variants have since been demonstrated.
@@ 3. Defining Security
To achieve security, we rely on both the hardware to enforce some isolation and on the fact the software verifies some requirements: we do not protect all-software out-of-the-box.
Intuitively, for a given security property, if the isolation enforced by the hardware is stronger, then the requirements on the software are weaker and we can enforce the security property for more programs.
Previous Work Previous work looking at enclaves in a context of transient execution attacks (bourgeat2019mi6, ) aims at limiting the attacker’s view to the enclave final output. Completion time was considered outside of the threat model and could be independently handled using traffic shaping techniques (padding) techniques. Unfortunately, this security property is too restrictive as it only considers a limited class of programs that do not communicate with the outside world.
Public Shared Memory We define a weaker security property in order to consider a more important class of programs to run securely. Our goal is for the system to guarantee that an attacker should only be able to observe the transcript of non-speculative accesses to a shared memory region. In particular, we don’t want an attacker to be able to observe any speculative accesses to shared memory from the victim nor any other microarchitectural state (accesses to private memory, branch predictors, TLBs etc…). However, similarly to previous work, we do not consider the timing of each access is observable by the attacker, i.e., there are no timestamps for the non-speculative shared memory accesses. We consider that leakage through the timing of shared memory accesses should be addressed using orthogonal traffic-shaping techniques such as (deutsch2022dagguise, ; wang2014timing, ; shafiee2015avoiding, ; zhou2017camouflage, ) and is left for future work.
Hardware & Software Requirements We now define what kind of requirements the hardware will needs to enforce and how the hardware restricts the victim program speculation model to help it achieve the just described security guarantee. 1) the hardware should properly isolate private microarchitectural state. 2) together, the hardware and the software need to prevent secrets from leaking through speculative accesses to shared memory. Intuitively, the more restrictive the speculation model of shared memory accesses enforced by the hardware, the lighter the conditions on the program to finally achieve security, but the least performance we get.
We first experiment with a restrictive speculative model where the victim program is not able to perform speculative accesses to shared memory. This effectively prevents the existence of any speculative transmitters between the victim and the attacker. The program only needs to make sure that the explicit shared memory accesses don’t leak information about the program secrets. Any constant-time cryptographic code satisfies these restrictions. Unfortunately this restrictive model gives us poor performance.
We experiment with a second speculative model called Burst mode where speculation on shared memory accesses is allowed, but all the predictors are deactivated and hard-coded to $pc+4$. The rationale is to allow pipelined accesses to shared memory to improve performance. Burst mode is intended to be only activated for code segments with straight-lined, burst accesses to shared memory. Under this less conservative speculation model, the software requirements are more important and the programs needs to verify that straight-line speculation ($pc+4$) within these segments will not leak any secrets. Outside of the said segments, we use the first speculation model.
@@ 3.1. Attacker Capabilities
We now complete our threat model by describing the attacker’s capabilities. The attacker is collocated on the same machine as the victim and can mount any software-based attack, including microarchitectural side channels and transient execution attacks. The attacker has compromised the majority of the software stack, including the operating system (OS) except for a limited trusted code base (TCB) including the secure bootloader, the SM (together 15K lines of code (LOC)) and the enclave. For the rest of this paper, we refer to any code outside of the TCB as the OS. We ignore attacks requiring physical access to the machine such as physical (e.g., noise, electromagnetic field, etc.) side channels. We also ignore side channels that can be mounted in software, but exploit the physical layer like power side channels. We do not consider Rowhammer attacks. As highlighted by our security definition, we are worried about enclave privacy and do not consider denial of service.
4. Citadel 1.0: A Secure Baseline
Even though defense mechanisms often propose to ”not speculate” when executing some instructions that can dirty the microarchitectural state, implementing such policies is not completely straightforward. In this section, we describe how to securely enable shared memory for our specific out-of-order open-source design while disabling speculative shared accesses. We also make sure to not enable any new speculative transmitters by bypassing the L1 D cache, and to not break address translation by tagging the TLB’s and the Address Translation Cache’s entries. Finally, we give a brief overview of the software infrastructure we have implemented in order to boot Linux and run enclaves from user programs.
4.1. Hardware Isolation

### v4

# Citadel: Real-World Hardware-Software Contracts for Secure Enclaves Through Microarchitectural Isolation and Controlled Speculation

@@ Abstract.
Hardware isolation primitives such as secure enclaves aim to protect sensitive programs, but remain vulnerable to transient execution attacks. Complete microarchitectural isolation is not a satisfactory defense mechanism as it leaves out public shared memory, critical for usability and application performance. Conversely, hardware-software co-designs for secure speculation can counter these attacks but are not yet practical, since they make assumptions on the speculation modes, the exposed microarchitectural state, and the software, which are all hard to support for the entire software stack. This paper advocates for processors to incorporate microarchitectural isolation primitives and mechanisms for controlled speculation, enabling different execution modes. These modes can restrict what is exposed to an attacker, effectively balancing performance and program-analysis complexity. We introduce two mechanisms to securely share memory between an enclave and an untrusted OS in an out-of-order processor. We show that our two modes are complementary, achieving speculative non-interference with a reasonable performance impact, while requiring minimal code annotation and simple program analysis doable by hand. Our prototype, Citadel, is a multicore processor running on an FPGA, booting untrusted Linux, and supporting comprehensive enclave capabilities, such as shared memory, and remote attestation. To our knowledge, Citadel is the first end-to-end enclave platform to run secure applications, such as cryptographic libraries or small private inference workloads, on a speculative out-of-order multicore processor while protecting against a significant class of side-channel attacks.
@@ 1. Introduction
Hardware isolation primitives such as Secure Enclaves or Trusted Execution Environments (TEEs) (costan2016sanctum, ; aegis, ; ferraiuolo2017komodo, ; lee2020keystone, ; bahmani2021cure, ; feng2021scalable, ) are used to run sensitive computations in isolation with strengthen privacy and integrity properties, even in the presence of a malicious Operating System (OS) or hypervisor (see Figure 1). Academic TEE proposals date back twenty years to XOM (xom, ) and AEGIS (aegis, ) and industry implementations include ARM TrustZone (trustzone04, ; pinto2019demystifying, ), Intel SGX (mckeen2013sgx, ; costan2016intel, ) and AMD SEV (sev2020strengthening, ). Unfortunately, these kinds of platforms have been shown to be vulnerable to microarchitectural side channels and transient execution attacks (chen2019sgxpectre, ; koruyeh2018spectre, ; ryan2019hardware, ; van2020lvi, ; van2018foreshadow, ; van2020sgaxe, ). One academic proposal (bourgeat2019mi6, ) has looked at securing enclaves in the context of microarchitectural side channels and transient execution attacks. They aim for total microarchitectural isolation between a (victim) enclave and the other security domains collocated on the same machine, leaving out shared memory, a commonplace mechanism at the center of many software and applications.
@@
On the other hand, mechanisms for secure speculation (ainsworth2020muontrap, ; cheang2019formal, ; canella2019systematic, ; barber2019specshield, ; fustos2019spectreguard, ; yan2018invisispec, ; yu2019speculative, ; saileshwar2019cleanupspec, ; weisse2019nda, ; khasawneh2019safespec, ; guarnieri2020spectector, ; guarnieri2021hardware, ; mosier2022axiomatic, ) can thwart these types of attacks while enabling shared resources but are not yet practical. Solutions that are purely hardware are complex: none has been implemented beyond simulators, and attacks have been found for the majority of schemes. Hardware-software co-design approaches simplify that problem by only providing guarantees for some programs, but they make strong assumptions on the type of speculative execution, the exposed microarchitectural state, and the software, which are hard to support for the entire software stack.
We propose that processors expose microarchitectural isolation primitives and mechanisms for controlled speculation. Indeed, we believe these two approaches go hand-in-hand as isolation makes it possible to constrain the problem of secure-speculation enough so hardware-software co-design solutions only require minimum hardware changes and a tractable program analysis component. To demonstrate this idea, we present two mechanisms that securely enable shared memory between an enclave and an untrusted OS on an out-of-order processor. Together, the enclave program and the hardware achieve Speculative Non-Interference (SNI) (guarnieri2020spectector, ; guarnieri2021hardware, ; behnia2021speculative, ; fabian2022automatic, ; yu2019speculative, ; yu2019speculative2, ) i.e., a victim program running in an enclave does not leak more secret information when running speculatively using our mechanisms than when running non-speculatively (one-instruction-at-a-time and in-order). For the rest of this paper, we will refer to non-speculative execution as sequential execution. Our security argument leverages hardware software contracts for security (guarnieri2021hardware, ; mosier2022axiomatic, ). Through microarchitectural isolation, the attacker observations are restricted to the addresses of victim accesses to shared memory. The hardware can then enforce some controlled speculative execution mode for the enclaves, such that an analyzed program can be run securely with regard to that mode. Intuitively, the more restrictive the execution mode, the lighter the requirements on the program, but the least performance we get.
Concretely, we first propose and build SpecSafe, a hardware mechanism that enforces an execution mode for the enclave where most instructions can be run speculatively, but no speculative shared accesses are allowed. This effectively prevents the existence of any speculative transmitters between the victim and the attacker without requiring any program modification or analysis. Unfortunately, we show it has limited performance on code that perform back-to-back accesses to shared memory. As a result, we present Burst Mode, a second enclave execution mode. We introduce software controls to enable Burst Mode where pipelined accesses to shared memory are allowed, but all the predictors of the executing core are hardwired to $pc+4$. This speculation model is usually referred to as “straight-line speculation” (straightlinespecARM, ) and we explain under which conditions a program can use it securely. We show that program analysis is tractable by hand on small straight-line code segments with back-to-back accesses to shared memory (e.g., a memcopy function). Our two execution modes are complementary, achieving SNI with a reasonable performance impact, while requiring minimal code annotation and simple program analysis doable by hand. To our knowledge, we are the first to prototype this type of security contract on a full platform, i.e., implemented on a processor running on FPGA booting Linux.
While building this prototype, we uncovered a few implementation subtleties and potential security issues that had been overlooked in the past, and we prototyped and run a simple-to-implement and effective Last Level Cache (LLC) isolation technique. Those side quests, though more implementation-centered, would be especially important to potential industrial deployment.
@@
Our final prototype is a modified out-of-order RISC-V 2-core processor synthesized to run on an AWS F1 FPGA. It fully implements all our security and performance mechanisms at the microarchitectural level, along with an end-to-end attestation mechanism. It boots (untrusted) Linux and user-level applications can safely launch and communicate with secure enclaves, executing applications such as a lightweight Python runtime, a cryptographic library or a private inference application running a small neural network. We implement and make all the pieces of software and modified hardware available, such as our modified processor, our secure bootloader, our security monitor and a Linux kernel module to interact with the security monitor. We summarize the different elements implemented for this paper in Figure 1. Citadel offers a unique environment for testing complex hardware-software mechanisms, making it ideal for research into microarchitectural side channels, transient execution attacks, and secure enclaves.
@@
  • To our knowledge, we show the first implementation leveraging hardware-software contracts to guarantee well-defined security properties.
  • We demonstrate how microarchitectural isolation goes hand-in-hand with controlled-speculation by presenting two mechanisms for public shared memory on enclaves running speculatively.
  • To our knowledge, we show the first end-to-end enclave platform prototype running securely on an FPGA implementation of a speculative out-of-order machine.
  • We show a hardware-software mechanism for cache partitioning that makes the platform more configurable and improves both enclaves’ and unprotected processes’ performance.
  • Through our case study, we provide important takeaways for commercial enclave offerings.
@@
2. Background
@@ 2.1. Security Monitor
In many enclave platforms, including Citadel, the Security Monitor (SM) (costan2016sanctum, ) (lebedev2019sanctorum, ) is a small (9K LOC in our prototype), trusted piece of software running at a higher privilege mode than the hypervisor or the OS (see Figure 1). Its role is to link low-level invariants exposed by the hardware (e.g., if a specific memory access is authorized), invoking the right microarchitectural cleanup routine, and the high-level security policies defined by the platform (e.g., which enclave is currently executing on which core and which memory region does it own).
@@ 2.2. Microarchitectural Timing Side Channels
To summarize the observations that could be made through side-channels, the community formalizes side-channels as a set of observations that an attacker can make about the microarchitectural state. For example, cache side-channels on unprotected machines typically lead to an observation mode that includes the sequence of addresses of memory operations (sometimes with their timestamps).
2.3. Transient Execution Attacks
Speculative or transient execution is one of the key performance features of modern processors. Unfortunately, it can make reasoning about security boundaries more difficult. Some microarchitectural data structures that orchestrate speculation (e.g., branch predictors) can be trained so the processor will execute code snippets speculatively, in ways that are incorrect or on data that should not be accessed during normal execution. Concretely, instructions that were not part of the normal program execution path can be executed out-of-thin-air, speculatively, by the processor. An example is the execution of code speculatively before the resolution of a preceding security check, such as a memory access bound. These erroneous executions happen speculatively, and these instructions will be squashed by the processor. Nonetheless, they will still change the microarchitectural state of the machine. For example, a speculative load to some data will not affect the values of the architectural registers, but will load the data in the memory hierarchy and possibly change the state of caches. This means that a program that manages sensitive data can be coerced into leaking private data speculatively through a side channel. More importantly, once some specific code snippet, or gadget, has been identified, these attacks can be used to leak virtually any data located in the victim’s virtual address space. This type of attack, called Spectre, was first introduced in (Kocher2018spectre, ), and many variants have since been demonstrated.
2.4. Hardware Software Contract for Security
A recent line of work (guarnieri2021hardware, ; mosier2022axiomatic, ) has been trying to define hardware software contracts for security, also called security contracts. As said in (guarnieri2021hardware, ), “the purpose of a contract is to split the responsibilities for preventing side channels between software and hardware”. An initial observation is that not all programs pose the same level of problem or difficulty in mitigation for all microarchitectural side-channels. This line of work tries to identify and describe classes of programs for which it is possible to implement performant microarchitecture securely, i.e., the programs should not leak more secrets through the microarchitecture implementation, than what it would directly leak in a non-speculative or sequential (one-instruction-at-a-time and in-order) Instruction Set Architecture model. In this work, we reuse the framework from (guarnieri2021hardware, ) while abstracting away some of its formalism. We point out that if we make some simple assumptions (that we check) on the parts of the enclave programs that are in charge of interacting with shared memory, we can run those parts of the enclave in a controlled speculation mode, preserving performance while ensuring strong security guarantees.
@@ 3. Defining Security
Previous Work Previous work looking at enclaves in a context of transient execution attacks (bourgeat2019mi6, ) enforces Strong Microarchitectural Isolation (SMI) by completely isolating security domains at the microarchitectural level. See Section 4.1 on hardware isolation. Unfortunately, this isolation forbids any shared architectural state between an enclave and the OS, disallowing critical functionalities such as shared memory. Any interaction between the enclave and the OS has to go through the SM and requires several (expensive) context switches, microarchitectural flushes and data copies from one domain to another. In (bourgeat2019mi6, ) enclave completion time is considered outside of the threat model. Generally, completion time can and should be handled independently using traffic shaping (padding) techniques.
Secure Shared Memory We relax SMI to allow some shared memory between the enclave and the OS. Nevertheless, to prevent transient execution attacks, we require that the enclave software and the hardware together achieve Speculative Non-Interference (SNI) (guarnieri2020spectector, ; guarnieri2021hardware, ; behnia2021speculative, ; fabian2022automatic, ; yu2019speculative, ; yu2019speculative2, ), i.e., that a victim program running in an enclave should not leak more secrets when running speculatively using our mechanisms than if run with non-speculatively (sequential execution). However, similar to previous work, we do not consider the timing of microarchitectural events to be visible to the attacker. In summary, no timestamps are associated with the observed shared memory accesses and the associated microarchitectural events; only the traces of the corresponding addresses are exposed. We consider that leakage through the timing of shared memory accesses is similar to leakage through completion time, and it should be addressed using orthogonal traffic-shaping techniques such as (deutsch2022dagguise, ; wang2014timing, ; shafiee2015avoiding, ; zhou2017camouflage, ). This is left for future work.
@@ 3.1. Attacker Capabilities
We now complete our threat model by describing the attacker’s capabilities. The attacker is collocated on the same machine as the victim enclave and can mount any software-based attack, including microarchitectural side channel and transient execution attacks. The attacker has compromised the majority of the software stack, including other malicious enclaves and the operating system (OS), except for a limited trusted code base (TCB) including the secure bootloader, the SM (together 15K lines of code (LOC)) and the victim program running inside of the enclave. The attacker also has control over the contents of shared memory. For the rest of this paper, we refer to any code outside of the TCB as the OS. We ignore attacks requiring physical access to the machine such as physical (e.g., noise, electromagnetic field, etc.) side channels. We also ignore side channels that can be mounted in software, but exploit the physical layer like power side channels. We do not consider Rowhammer attacks. As highlighted by our security definition, we are worried about enclave privacy and do not consider denial of service. We also do not protect enclave programs that explicitly leak their own secrets, including through shared memory.
4. Secure Shared Memory
In this section, we first describe our baseline microarchitecture, and how it enforces strong isolation between security domains with shared memory disabled. We then describe how to enable shared memory, how it opens the door to Spectre-like attacks, and why existing mechanisms are insufficient to address them. We then describe SafeSpec and Burst Mode, our two complementary mechanisms to securely and efficiently enable shared memory on speculative microarchitectures.
4.1. Strong Microarchitectural Isolation
Figure 2. Resource Partitioning Between Security Domains
Figure 3. Context Switching Between Security Domains